OpenStudy (anonymous):

Whats the radius of a 4 mm^2 electrical copper wire?

OpenStudy (anonymous):

1.1286529

OpenStudy (anonymous):

Can you explain the process finding that answer?

OpenStudy (anonymous):

u r giving the square lenth mm^2 means cross sectional area ryt and that is pi*r^2=3.14*r^2=4... now r^2=4/3.14.. u can calculate r from here

OpenStudy (anonymous):

Okey, thanks!

OpenStudy (anonymous):

most welkum